OpenStudy (anonymous):

Let the function f be defined by y = f(x), where x and f(x) are real numbers. Find the following. (Simplify your answers completely.) f(x)=3/7x+5 f(2) = f(−3) = f(k) = f(k2 − 1) =

OpenStudy (anonymous):

f(2) = 3/7(2)+5 =3/14+5 =73/14 similarly f(-3) = 3/7(-3)+5 =3/-21 +5 =-102/-21 =102/21 f(k) = (3+35k)/7 f(k\(^{2}\)) = (3+35k\(^{2}\))/7

OpenStudy (anonymous):

is this \[f(x)=\frac{3}{7x+5}\]

OpenStudy (anonymous):

if so then \[f(2)=\frac{3}{7\times 2+5}=\frac{3}{19}\]

OpenStudy (anonymous):

and likewise \[f(-3)=\frac{3}{7\times -3+5}=-\frac{3}{16}\]\]